OpenStudy (anonymous):

Ok, so i know that C = 2"pi"r so that meant that pi= c/2r?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

"means"

OpenStudy (anonymous):

\[C=2\pi r\]\[{C\over2}=\pi r\]\[{C\over{2r}}=\pi\]\[\large\pi={C\over2r}\]

OpenStudy (anonymous):

Ok! Thanks so much for checking that for me!@yummydum

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks :) @radar

OpenStudy (radar):

You're welcome and your reasoning is on track.

OpenStudy (anonymous):

Ok :) Thank you!